What is the specific volume of air at 15 psia and 90F in ft^3/lb?
1
Expert's answer
2021-10-06T15:52:56-0400

T=90 F = 549.67 °R

"p=15 \times 144 = 2160 \;lb/ft^2"

"R=1.554 \times 10^3"

Ideal gas law

"pV=RT \\

2160V = \frac{1.554 \times 10^3}{28.67} \times 549.67 \\

V=13.79 \;ft^3/lb"